# Spatial-Aware Conformal Prediction for Trustworthy Hyperspectral Image Classification

### v1

@@ V-C Discussion
SACP Exhibits Adaptiveness
Adaptiveness is a crucial attribute of CP, ensuring that prediction sets accurately reflect the uncertainty associated with each instance. Specifically, prediction sets should be larger for more challenging instances and smaller for easier ones. This analysis evaluates whether SACP maintains this characteristic. We use the true label rank $o(y,\hat{\pi}_{\theta}(\bm{x}))$ to represent the difficulty of the instance $\bm{x}$, where a higher true label rank indicates greater classification difficulty. Instances are categorized into five groups based on their true label rank: $1$, $2$, $3$, $4$, and $\{5,\cdots,|\mathcal{Y}|\}$. We then compute the average size of prediction sets for each group. This experiment is conducted on the SA and PU datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$. Figure 4 demonstrates that SACP exhibits adaptiveness across various datasets. Both APS, with or without SACP, consistently show that prediction sets are smaller for easier instances compared to more challenging ones. SACP effectively optimizes the average size of prediction sets for each difficulty group while maintaining its adaptive characteristics. These results collectively underscore the significance of SACP in HSI classification.
@@
In this section, we analyze the computational efficiency of SACP. For the base score functions APS and RAPS, the computation involves operations such as sorting and summation. Consequently, the time complexity for computing both basic scores is $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N)$, where $N$ denotes the number of instances and $|\mathcal{Y}|$ represents the total number of labels. The dominant computational overhead in this process arises from the sorting operation. SACP introduces an additional aggregation step, contributing an extra time complexity of $\mathcal{O}(kN)$, where $k$ is the number of iterations. Thus, the overall time complexity of SACP is $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N+kN)$. For simplicity, we approximate $(k+|\mathcal{Y}|\log|\mathcal{Y}|)$ as a constant $C$, since these terms are negligible compared to $N$. Thus, the complexity can be expressed as $\mathcal{O}(CN)$, which remains comparable in magnitude to that of standard CP. The above analysis demonstrates that SACP exhibits computational efficiency.
@@ VI Conclusion
In this paper, we investigate the application of CP to HSI classification and introduce a conformal procedure that integrates conformal prediction sets into HSI classifiers, which ensure the inclusion of true labels with a user-specified probability. Building on this, we propose SACP, a framework designed to enhance the efficiency of CP for HSI classification. The key idea of SACP is its utilization of intrinsic spatial information in HSIs to refine non-conformity scores. To the best of our knowledge, this is the first study to explore CP within the context of HSI data. Extensive experiments reveal that SACP outperforms standard CP in HSI classification by reducing the average prediction set size while ensuring satisfactory marginal and conditional coverage rates. Moreover, SACP is practical and user-friendly, showing robustness to hyperparameters and incurring minimal computational cost.
@@ Appendix A Proof of Theorem 2
As a result, any score of test sample in $\{s_{n+j}\}_{j=1}^{m}$ is exchangeable with $\{s_{i}\}_{i=1}^{n}$. Then, we complete the proof. ∎
@@ Appendix B Proof of Proposition 1
Moreover, the event $\{S(B,y)\leq\tau^{\alpha}\}$ in Eq. (23) is equivalent to the event that at least $\alpha$ percent of the sample in $\mathcal{D}_{\text{cal}}$ have scores larger than that of $(B,y)$. . Thus, we have

### v2

@@ V-B Main Results
SACP is insensitive to hyperparameters
In SACP, two hyperparameters are crucial: $\lambda$ and $k$. The parameter $\lambda$ serves as a trade-off, balancing the influence of the neighborhood, while $k$ specifies the number of iterations. Increasing $k$ allows SACP to incorporate spatial information from progressively larger neighborhoods. This section examines the impact of variations in $\lambda$ and $k$ on the performance of SACP. The experiment is conducted on all selected datasets using SSTN, with APS as the score function and $\alpha$ set to 0.05. The results are illustrated in Figure 3. Figure 3a demonstrates that SACP consistently outperforms the baseline across different values of $\lambda$, with the size of prediction sets decreasing as $\lambda$ increases. Figure 3b illustrates that SACP consistently achieves superior performance as the number of iterations $k$ changes. Overall, the results suggest that SACP is relatively insensitive to variations in $\lambda$ and $k$, consistently surpassing the baseline. For simplicity, we employ a moderate range of hyperparameters, consistently setting $\lambda=0.5$ and $k=1$ throughout the experiments. Further tuning these parameters on a hold-out set may further enhance the performance of conformal prediction.
SACP is robust to the size of calibration set
As shown in Eq. (2), the calibration set is used to determine the threshold $\tau$ corresponding to the coverage rate of $1-\alpha$. A precise estimate of $\tau$ is essential. In this work, all experiments are conducted with a fixed ratio for the calibration set size, specifically $|\mathcal{D}_{\text{cal}}|\approx 0.5\times|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$. In this section, we investigate whether SACP is sensitive to the size of the calibration set $|\mathcal{D}_{\text{cal}}|$. Specifically, we exam whether Coverage and Size remain satisfactory when varying the ratio of the calibration set size. The experiment is conducted on all three datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$. The ratio of calibration samples is defined as $\gamma=(|\mathcal{D}_{\text{cal}}|/|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|)$. $\gamma$ ranges from 0.1 to 0.6 in increments of 0.05. The experimental results are presented in Table IV. Coverage consistently maintains $1-\alpha$ as the calibration set size decreases relative to $|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$, even for very small values of $\gamma$. Additionally, the size of the prediction sets remains stable, only marginally influenced by changes in the calibration set size, and fluctuates within a narrow range. These findings collectively demonstrate that SACP exhibits significant robustness to variations in the calibration set size.
V-C Discussion
SACP exhibits adaptiveness
Adaptiveness is a crucial attribute of conformal prediction, ensuring that prediction sets accurately reflect the uncertainty associated with each instance. Specifically, prediction sets should be larger for more challenging instances and smaller for easier ones. This analysis evaluates whether SACP maintains this characteristic. We use the true label rank $o(y,\hat{\pi}_{\theta}(\bm{x}))$ to represent the difficulty of the instance $\bm{x}$, where a higher true label rank indicates greater classification difficulty. Instances are categorized into five groups based on their true label rank: $1$, $2$, $3$, $4$, and $\{5,\ldots,|\mathcal{Y}|\}$. We then compute the average size of prediction sets for each group. This experiment is conducted on the SA and PU datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$.
Figure 4 demonstrates that SACP exhibits adaptiveness across various datasets. In both cases, with or without SACP, APS consistently shows that prediction sets are smaller for easier instances compared to more challenging ones. We conclude that SACP effectively optimizes the average prediction set size for each difficulty group while preserving its adaptive characteristics. These results collectively underscore the significance of SACP in HSI classification.
@@ V-C Discussion
In this section, we analyze the computational efficiency of SACP. The base non-conformity score functions APS, RAPS, and SAPS involve only sorting and summation operations, resulting in a time complexity of $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N)$, where $N$ is the number of instances and $|\mathcal{Y}|$ represents the total number of labels. The dominant computational overhead arises from the sorting operation. In comparison, SACP introduces an additional aggregation step, which adds a time complexity of $\mathcal{O}(kN)$, where $k$ is the number of iterations. Consequently, the overall time complexity of SACP is $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N+kN)$. For simplicity, we approximate $(k+|\mathcal{Y}|\log|\mathcal{Y}|)$ as a constant $C$, as these terms are negligible compared to $N$. Therefore, the complexity can be expressed as $\mathcal{O}(CN)$, which is comparable in magnitude to that of standard conformal prediction. This analysis demonstrates that SACP exhibits computational efficiency.
@@ VI Conclusion
In this paper, we first present a comprehensive theoretical proof establishing the validity of conformal prediction for HSI classification. We then introduce a conformal procedure that integrates conformal prediction sets into HSI classifiers, ensuring the inclusion of true labels with a user-specified probability. Building on this foundation, we propose SACP, a conformal prediction framework specifically designed for HSI data. SACP refines non-conformity scores by leveraging the intrinsic spatial information present in HSIs, effectively enhancing the efficiency of conformal prediction sets. To the best of our knowledge, this is the first study to explore conformal prediction in the context of HSI data. Extensive experiments confirm the validity of conformal prediction for HSI classification. Moreover, SACP outperforms standard conformal prediction in HSI classification by reducing the average prediction set size while maintaining satisfactory marginal and conditional coverage rates. Furthermore, SACP is practical and user-friendly, exhibiting robustness to hyperparameter variations and incurring minimal computational costs.
@@ Appendix A Proof of Theorem 2
As a result, any score of test sample in $\{s_{n+j}\}_{j=1}^{m}$ is exchangeable with $\{s_{i}\}_{i=1}^{n}$. ∎
@@ Appendix B Proof of Proposition 1
Moreover, the event $\{S(B,y)\leq\tau^{\alpha}\}$ in Eq. (12) is equivalent to the event that at least $\alpha$ percent of the sample in $\mathcal{D}_{\text{cal}}$ have scores larger than that of $(B,y)$. Thus, we have